Question5 3 pts A woman drives a car from one city to another with different constant speeds along the trip. She drives at a speed of 26 km/h for 2.0 min, 58 km/h for 15.0 min, makes a stop for 41 min, then continues at 26 km/h for 44 min, at which point she reaches her destination. What is her average velocity for the entire trip, in km/hr?
Answer #1

distance = speed * time

d1 = 26 * (20/60) = 8.67 km

d2 = 58 * ( 15/60) = 14.5 km

d3 = 26 * (44/60) = 19.07 km

av velocity,

Vav = (d1 + d2+d3) / ( t1 + t2 + t3 + t4)

Vav = (8.67 + 14.5 + 19.07)* 60/ ( 20+ 15 + 41 + 44)

Vav = 21.12 km/h

========

Comment in case any doubt, will reply for sure.. Goodluck